OpenStudy (anonymous):

find if the seq converges or diverges {(-1)^n (n+5/5^n)}

OpenStudy (anonymous):

let \( a_n \) be the n-th element of the sequence. show that \[ i) \lim_{n \rightarrow \infty} a_n = 0 \]

OpenStudy (anonymous):

i found first term -6/5 second term -23/25 and the third term as -123/125 but i don't any relationship btw them

OpenStudy (anonymous):

you don't need relationship between them to show that the sequence converges. just show that as \( n \rightarrow \infty \) the value if the n-th term is zero. In other words, show that \[ \lim_{n\rightarrow \infty} {(-1)^n (n+5) \over 5^n} = 0\]

OpenStudy (anonymous):

If, \[ \lim_{n\rightarrow \infty} {(-1)^n (n+5) \over 5^n} = \infty (\text{ the series diverges} )\] \[ \lim_{n\rightarrow \infty} {(-1)^n (n+5) \over 5^n} = a \text{ (for some value 'a', then the series converges})\]

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

sorry, *sequence .. not series

OpenStudy (anonymous):

with the limit to infinity can you take out (1/5)^n

OpenStudy (anonymous):

use L'Hopital rule

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do u d/dx 5^n

OpenStudy (anonymous):

or (-1)^n

OpenStudy (anonymous):

your variable is 'n' not 'x' \[ {d \over dn} (5^n) = {d \over dn} (e^{n\ln 5})\] you should be able to do it. for (-1)^n, don't bother, just ignore it. If you want to do it then use power rule.